You plan to invest $25,000 in two funds paying 3% and 4 1/2% simple interest.(There is more risk in the 4 1/2% fund.) Your goal is to obtain a total annual interest income of $1000 from the investments. What is the smallest you can invest in the 4 1/2% fund in order to meet your objective?
Answer by stanbon(75887)

You plan to invest $25,000 in two funds paying 3% and 4 1/2% simple interest.(There is more risk in the 4 1/2% fund.) Your goal is to obtain a total annual interest income of $1000 from the investments. What is the smallest you can invest in the 4 1/2% fund in order to meet your objective?
-----
Equations:
Quantity: x + y = 25000
Interest:0.03x + 0.045y = 1000
-----------------------------------
Multiply thru 1st by 3
Multiply thru 2nd by 100
---------------------------
3x + 3y = 3*25000
3x+4.5y = 100,000
-------------------------
1.5y = 25000
y = 1666.67 (amt to invest at 4.5%
====================================
Cheers,
Stan H.